In the Haemoglobin (Molecular wt = 67200) iron found 0.33% (by weight). The number of iron atom will be in its one molecule 
  1. 1
  2. 2
  3. 3
  4. 4
Urgent answer required

Grade:12th pass

1 Answers

Bhavya
8 years ago
Since iron is found 0.33% by weight
it means in 100 g of Hb. iron present = 0.33
in 67200 g of Hb, iron present = (0.33/100)*67200
Now atleast 1 atom should be present in its 1 molecule = 56g
56 g = 1 atom
(0.33/100)*67200 = (1/56)*(0.33/100)*67200
= 4